The proportion of vehicles which drive above the speed limit on a freeway is 78%. Suppose 10 vehicles are randomly clocked. The following is the probability distribution of the number of vehicles which drive above the speed limit. (The blank cells are intentional!)
Answer #1

a) P(X=x)=n_{C_{x}}(1-p)^{n-x}(p)^{x}

P(X = 8) = 10G (1-078)10-8 (0.78) = 45(0.22)2(0.78) = 0.2984 C

b) P(X < 8)-1-P(X 9) _ P(X = 10) 1-0.2351-0.0834-0.6815 C